Day Shiftsum 15-Jan-2000


Subject: Day Shiftsum 15-Jan-2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:25:53 CST


        Shift Summary, Day shift, 15/Jan/2000

On Shift: Elizabeth T., Albert E.

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.5
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 4.25

Accounting for lost time:
Bad tape or tape unit probably. Changing over to KTeV99 configuration.
Linac vacuum problem.

========================================================================
Runs ended this shift: 15533, 15536,15537,15538,15539
Run 15540 BEAM_799 in progress

Tapes brought to FCC this shift:
QKE155-172, QKB165-166
========================================================================
Accesses during shift: One, to restore setup of E799

========================================================================
Overview of shift:

When tapes were changed for a new run by the previous shift, DAQ would
take data for several spills and stop. After several more spills it
would begin again for 3-5 more spills logging data. Ed came in a decided
to reboot KTeV1, replace all tapes, clean tape units, recycle tape power.
This produced smooth running until shutdown at 10;30 to switch to KTeV99.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
             7:00 DAQ takes data for several spills and rests
                        for several spills.
             7:50 Complete new set of tapes and an overhaul of
                        tape units restores peaceful running.
            10:30 Bob T. orders run stop of 15533 and a shift to
                        KTeV99 data taking with new tape set. Bob T.
                        successfully frees up stuck door on a tape unit
                        and puts it back in service
            12:20 Begin muon run 15536
            12:35 Jason L. does DC plateaus with runs
                        15537 and 15538
            13:32 Pedestal run 15539
            13:50 Begin KTeV99 run 15540. Linac develops vacuum
                        problems before run really gets going.
            15:00 Estimate beam at 15:15

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:25:55 CST